# Can LLMs Learn Macroeconomic Narratives from Social Media?

### v1

@@ Appendix E T5 Integration Prompt
Figure 7 shows a prompt example for the T5 model classifying ’direction change’ of the S&P 500 for 1-day prediction horizon. This model is used for the experiments in Section 7.2. The prompt contains both LLM-generated analysis and historical financial values, separated using leading titles and special tokens. The prompt ends with a title of the task itself. More details about the T and F integration can be seen in Section 4.3.
@@
Figure 7: Prompt example of the T5 prediction model receiving both LLM-generated analysis and historical financial values, in order to classify ’direction change’ of the S&P 500.
@@ F.1 Twitter Analysis
To inspect the information contained in our Pre-Pandemic Twitter dataset, we conduct an analysis of the 100 most frequent words associated with each topic of the dataset. This analysis reveals temporal shifts in the distribution of most frequent words, coinciding with significant world events. For instance, Figure 8 illustrates the monthly number of occurrences of the term ’debt’ in the ’economics’ dataset, presenting peaks in July 2011 and October 2013, corresponding to major debt ceiling crises experienced by the US during these times.
Figure 9 shows (a) sentiment scores created by VADER (Hutto and Gilbert, 2014) sentiment model (ranged [-1,1]) for the ’business’ related tweets; and (b) the sentiment scores divided to bins of size 0.5.
Figure 8: Frequency of the word ’debt’ in the ’economics’ dataset over the years. Observed peaks in July 2011 and October 2013 are aligned with the major debt ceiling crises the US experienced in these times.
@@
Figure 9: VADER (Hutto and Gilbert, 2014) sentiment scores of the ’business’ dataset.
@@ F.2 LLM-Based Narrative Analyses Examples
Appendix G Full Results for ’Sentiment-Based Next-Day Prediction’ Experiment
This appendix shows the full results for the experiment described in Section 6.1. Tables 6 and 7 presents the performance of all compared models for both ’direction change’ and ’next value’ tasks. For the summary of results and takeaways, please refer to the main experiment section.
@@ Appendix G Full Results for ’Sentiment-Based Next-Day Prediction’ Experiment
Table 6: ’Direction change’ classification of FFR. Financial features are encoded either as ’Binary’ (increase/decrease) or ’Numeric’. The TF and F models perform nearly identically, outperforming the T models and the F baselines. Best performing models are F & TF Random Forest with numeric financial features.
@@
Table 7: ’Next value’ prediction of FFR. Best T, F, TF models are comparable to each other and to ’train-mean’ F baseline.
Appendix H Full Description for ’LLMs for End-to-End Prediction’ Experiment
@@
Synthetic Scenarios: We manipulate input data to investigate the LLM’s grasp of causal relationships, feeding it with texts coupled with corresponding financial changes (see the ’synthetic narratives’ baseline in § 3). While the LLM recognizes the direction of impact of these scenarios, it struggles to quantify the actual magnitude of the change.

### v2

@@ Appendix E T5 Integration Prompt
Figure 7 shows a prompt example for the T5 model classifying “direction change” of the S&P 500 for 1-day prediction horizon. This model is used for the experiments in Section 8.2. The prompt contains both LLM-generated analysis and historical financial values, separated using leading titles and special tokens. The prompt ends with a title of the task itself. More details about the T and F integration can be seen in Section 5.3.
@@
Figure 7: Prompt example of the T5 prediction model receiving both LLM-generated analysis and historical financial values, in order to classify “direction change” of the S&P 500.
@@ F.1 Twitter Analysis
To inspect the information contained in our Pre-Pandemic Twitter dataset, we conduct an analysis of the 100 most frequent words associated with each topic of the dataset. This analysis reveals temporal shifts in the distribution of most frequent words, coinciding with significant world events. For instance, Figure 8 illustrates the monthly number of occurrences of the term “debt” in the “economics” dataset, presenting peaks in July 2011 and October 2013, corresponding to major debt ceiling crises experienced by the US during these times.
Figure 9 shows (a) sentiment scores created by VADER (Hutto and Gilbert, 2014) sentiment model (ranged [-1,1]) for the “business” related tweets; and (b) the sentiment scores divided to bins of size 0.5.
Figure 8: Frequency of the word “debt” in the “economics” dataset over the years. Observed peaks in July 2011 and October 2013 are aligned with the major debt ceiling crises the US experienced in these times.
@@
Figure 9: VADER (Hutto and Gilbert, 2014) sentiment scores of the “business” dataset.
@@ F.2 LLM-Based Narrative Analyses Examples
Appendix G Full Results for “Sentiment-Based Next-Day Prediction” Experiment
This appendix shows the full results for the experiment described in Section 7.1. Tables 6 and 7 presents the performance of all compared models for both “direction change” and “next value” tasks. For the summary of results and takeaways, please refer to the main experiment section.
@@ Appendix G Full Results for “Sentiment-Based Next-Day Prediction” Experiment
Table 6: “Direction change” classification of FFR. Financial features are encoded either as “Binary” (increase/decrease) or “Numeric”. The TF and F models perform nearly identically, outperforming the T models and the F baselines. Best performing models are F & TF Random Forest with numeric financial features.
@@
Table 7: “Next value” prediction of FFR. Best T, F, TF models are comparable to each other and to “train-mean” F baseline.
Appendix H Full Description for “LLMs for End-to-End Prediction” Experiment
@@
Synthetic Scenarios: We manipulate input data to investigate the LLM’s grasp of causal relationships, feeding it with texts coupled with corresponding financial changes (see the “synthetic narratives” baseline in § 4). While the LLM recognizes the direction of impact of these scenarios, it struggles to quantify the actual magnitude of the change.